Two trains are lewaving a city on parallel tracks, travelling in the same direction. The passenger train is going twice as fast as the freight train. After 45 minutes, the trains are 180 miles apart. Find the speed of each train.
Answer by Alan3354(21603)

Two trains are lewaving a city on parallel tracks, travelling in the same direction. The passenger train is going twice as fast as the freight train. After 45 minutes, the trains are 180 miles apart. Find the speed of each train.
------------------------
180 miles in 45 minutes --> 240 mph
240 mph is the difference in the speeds, since they're going the same direction.
--------------------
r1 = r2+ 240
r1 = 2*r2
2*r2 = r2 + 240
r2 = 240 mph (speed of one train)
r1 = 480 mph (speed of 2nd train)
Those are some fast trains. I think the problem said different directions.
------------------------
Solution going opposite directions:
r1 + r2 = 240
r1 = 2*r2
----------
Sub r1 into 1st eqn
2*r2 + r2 = 240
3*r2 = 240
r2 = 80 mph
r1 = 160 mph
Still some fast movers, but more likely.